# Small Reward Models via Backward Inference

### v1

@@ 2.2 Our method: FLIP
Ablation studies on similarity metrics and robustness to prompt are presented in Appendix G, E.
@@ 3.3 Reinforcement Learning
We use the DPO-trained version of OLMo3-7B-Think (Olmo et al., 2025), which has not been trained with RLVR, as the base policy. We sample 12k English prompts from the WildChat (Zhao et al., 2024), featuring real-world ChatGPT conversations. For each prompt, we generate $K=8$ rollouts with a temperature of 1.0. We train with a response length of 16K tokens with learning rate of $1\times 10^{-6}$. Training is performed until convergence. We evaluate the effectiveness of FLIP against the Pointwise Rating variant, which is the only variant applicable in this setting, using Qwen3-1.7B and Qwen3-4B. Reward models are given a maximum context length of 25k tokens. Additional experimental details are in Appendix I.
@@ 4.1 Intrinsic Evaluation
from below to above random chance
@@
smaller the model, greater the boost
@@
on the Focus subset, FLIP using small models is comparable to LLM-as-a-Judge using large commercial models
@@
Method claude gpt
3-Haiku 3-Opus Sonnet-4 4.1-mini 4.1 4o
LLM-as-a-Judge 50.1 66.5 76 73.5 73.4 72.9
@@ 4.2 Test-Time Scaling with Parallel Sampling
Figure 4: Test-time scaling with parallel sampling results. OLMo 2 results are averaged across 1B and 7B instruct variants, while Llama 3 results are averaged across 1B, 3B, and 8B instruct variants. See results of individual models in Appendix B. FLIP substantially outperforms the baselines, achieving higher performance with greater stability.
@@ 5 Analysis
  • Off-topic or irrelevant responses. (Tables 14, 15) When a response deviates from the instruction, the inferred instruction will diverge significantly, resulting in very low similarity scores.
@@
  • Under- or over-addressed responses. (Tables 17, 18) If the response includes content that is not requested by the instruction, or addresses only a subset of the required components, the inferred instruction may introduce additional constraints or omit parts of the original instruction, leading to a noticeable mismatch.
@@
We further examine the results on RewardBench2 and analyze how response length affects the performance. Specifically, we select instances in which all four candidate responses fall into the same length category—short, medium, or long. This constitutes a challenging setting, as the models are tasked with differentiating between responses of similar lengths. The results are presented in Figure 5. Across all length groups, FLIP consistently outperforms LLM-as-a-Judge, with the most pronounced gains observed in the long response category. As response length increases, models are provided with richer contextual information for backward inference, and the mapping between responses and their underlying instructions becomes closer to one-to-one. Results with respect to generator model size are in Appendix D.
@@ 6 Related Work
Reward modeling in non-verifiable domains primarily relies on LLM-as-a-Judge (Zheng et al., 2023). However, a growing body of work demonstrates that LLM judges exhibit systematic biases (Zheng et al., 2023; Jung et al., 2025; Ye et al., 2024; Chen et al., 2024a; Li et al., 2025; Zeng et al., 2023; Zhao et al., 2025) and are vulnerable to adversarial attacks that artificially inflate evaluation scores through carefully crafted text sequences (Raina et al., 2024; Shi et al., 2024a). As alternatives, several methods depend on reference responses (Lyu et al., 2024). Chang et al. (2025) computes BLEU scores (Papineni et al., 2002) between sampled outputs and reference answers, while Yu et al. (2025a) leverages the token-level probability assigned to reference responses. Another line of research employs rubric-based evaluation (Gunjal et al., 2025; Viswanathan et al., 2025), decomposing desirable response qualities into human-interpretable criteria. Nevertheless, constructing high-quality reference answers or well-defined rubrics remains costly and challenging, particularly at scale. We instead propose FLIP, which operates effectively with small models in downscaled regimes without requiring references or evaluation rubrics.
@@ 7 Conclusion
First, one potential confounding factor for FLIP is that some responses may begin by fully or partially repeating the instruction. We offer three observations regarding this issue. First, based on our empirical analysis, such cases are relatively rare, although their frequency varies across models, and this behavior can be readily discouraged or detected using simple heuristics. Second, responses that begin by repeating the instruction are often high-quality and exhibit strong instruction-following behavior, which mitigates their confounding effect on reward estimation. Third, even when a response partially repeats the instruction, the reward model does not simply reproduce the repeated text as the inferred instruction; instead, the inferred instruction remains dependent on the substantive content of the response. Therefore, we argue that the overall impact of this issue on FLIP is limited.
Second, FLIP relies on the F1 score, the harmonic mean of word-level precision and recall, as the similarity metric between the inferred instruction and the original instruction. While effective in most settings, this metric becomes inappropriate in rare cases where the inferred instruction and the original instruction are expressed in different languages. In such scenarios, we recommend replacing the F1 score with a LLM judge (Appendix G).
@@ Appendix A Future Work
Appendix B Test-Time Scaling with Parallel Sampling
Figures 7 and 8 present detailed results of test-time scaling with parallel sampling using the OLMo 2 and Llama 3 model families as reward models, respectively.
Appendix C Qualitative Examples
Appendix D Generator Size
Appendix E Robustness to Prompt Variations
@@
Table 4: Accuracy of the Llama 3 family models, averaged across the 1B, 3B, and 8B instruction-tuned variants, using various prompts on RewardBench2 across all subsets. “s.d.” denotes the standard deviation across different 3 prompts for each method. See Appendix F for the full list of prompts. FLIP is more robust to prompt variations than LLM-as-a-Judge.
Appendix F Prompts
We list the prompts used in the main sets of experiments in Tables 5 and 6. Additional prompts used for the prompt consistency check are provided in Tables 7 through 10.
Table 5: Prompts used for all experiments, unless otherwise stated.
Table 6: Prompts used for all experiments, unless otherwise stated.
Appendix G Similarity Metric
@@
Appendix H Scaling Up
@@
Appendix I Experiment Details
@@
Inference experiments are performed on either A100 or H200 GPUs. If a response does not follow the instructed output format, it is counted as incorrect.

### v2

@@ 2.2 Our method: FLIP
Ablation studies on similarity metrics and robustness to prompt are presented in Appendix D, C.
@@ 3.3 Reinforcement Learning
We use the DPO-trained version of OLMo3-7B-Think (Olmo et al., 2025), which has not been trained with RLVR, as the base policy. We sample 12k English prompts from the WildChat (Zhao et al., 2024), featuring real-world ChatGPT conversations. For each prompt, we generate $K=8$ rollouts with a temperature of 1.0. We train with a response length of 16K tokens with learning rate of $1\times 10^{-6}$. Training is performed until convergence. We evaluate the effectiveness of FLIP against the Pointwise Rating variant, which is the only variant applicable in this setting, using Qwen3-1.7B and Qwen3-4B. Reward models are given a maximum context length of 25k tokens. Additional experimental details are in Appendix H.
@@ 4.1 Intrinsic Evaluation
FLIP outperforms LLM-as-a-Judge by 79.6% on small language models.
@@
Smaller the model, greater the boost.
@@
On the Focus subset, FLIP using small models is comparable to LLM-as-a-Judge using large commercial models.
@@
Method Claude-Sonnet-4 GPT-4.1-mini GPT-4.1
LLM-as-a-Judge 76 73.5 73.4
FLIP 77.7 75.2 75.0
@@ 4.2 Test-Time Scaling with Parallel Sampling
Figure 4: Test-time scaling with parallel sampling results. OLMo 2 results are averaged across 1B and 7B instruct variants, while Llama 3 results are averaged across 1B, 3B, and 8B instruct variants. See results of individual models in Appendix G. FLIP substantially outperforms the baselines, achieving higher performance with greater stability.
@@ 5 Analysis
  • Off-topic or irrelevant responses. (Tables 15, 16) When a response deviates from the instruction, the inferred instruction will diverge significantly, resulting in very low similarity scores.
@@
  • Under- or over-addressed responses. (Tables 18, 19) If the response includes content that is not requested by the instruction, or addresses only a subset of the required components, the inferred instruction may introduce additional constraints or omit parts of the original instruction, leading to a noticeable mismatch.
@@
We further examine the results on RewardBench2 and analyze how response length affects the performance. Specifically, we select instances in which all four candidate responses fall into the same length category—short, medium, or long. This constitutes a challenging setting, as the models are tasked with differentiating between responses of similar lengths. The results are presented in Figure 5. Across all length groups, FLIP consistently outperforms LLM-as-a-Judge, with the most pronounced gains observed in the long response category. As response length increases, models are provided with richer contextual information for backward inference, and the mapping between responses and their underlying instructions becomes closer to one-to-one. Results with respect to generator model size are in Appendix E.
@@ 6 Related Work
Reward modeling in non-verifiable domains primarily relies on LLM-as-a-Judge (Zheng et al., 2023). However, a growing body of work demonstrates that LLM judges exhibit systematic biases (Zheng et al., 2023; Jung et al., 2025; Ye et al., 2024; Chen et al., 2024a; Li et al., 2025; Zeng et al., 2023; Zhao et al., 2025) and are vulnerable to adversarial attacks that artificially inflate evaluation scores through carefully crafted text sequences (Raina et al., 2024; Shi et al., 2024a). As alternatives, several methods depend on reference responses (Lyu et al., 2024). Chang et al. (2025) computes BLEU scores (Papineni et al., 2002) between sampled outputs and reference answers, while Yu et al. (2025a) leverages the token-level probability assigned to reference responses. Another line of research employs rubric-based evaluation (Gunjal et al., 2025; Viswanathan et al., 2025; Dineen et al., 2025), decomposing desirable response qualities into human-interpretable criteria. Nevertheless, constructing high-quality reference answers or well-defined rubrics remains costly and challenging, particularly at scale. We instead propose FLIP, which operates effectively with small models in downscaled regimes without requiring references or evaluation rubrics.
@@ 7 Conclusion
FLIP relies on the premise that the response contains sufficient length and contextual richness. While this assumption holds in most cases, it does not apply to single-word or single-phrase responses, as noted in Footnote 3.
While, theoretically, multiple distinct instructions may yield the same response, we did not observe any instances in our dataset in which different questions shared an identical answer. For responses that incorporate adequate detail and contextual specificity, the corresponding set of plausible generating questions is typically severely restricted, approximating a one-to-one mapping between question and response.
The other potential confounding factor for FLIP is that some responses may begin by fully or partially repeating the instruction. We offer three observations regarding this issue. First, based on our empirical analysis, such cases are relatively rare, although their frequency varies across models, and this behavior can be readily discouraged or detected using simple heuristics. Second, responses that begin by repeating the instruction are often high-quality and exhibit strong instruction-following behavior, which mitigates their confounding effect on reward estimation. Third, even when a response partially repeats the instruction, the reward model does not simply reproduce the repeated text as the inferred instruction; instead, the inferred instruction remains dependent on the substantive content of the response. Therefore, we argue that the overall impact of this issue on FLIP is limited.
FLIP relies on word-level F1 score, the harmonic mean of word-level precision and recall, as the similarity metric between the inferred instruction and the original instruction. While effective in most settings, this metric becomes inappropriate in rare cases where the inferred instruction and the original instruction are expressed in different languages. In such scenarios, we recommend replacing the F1 score with a LLM judge (Appendix D).
@@ Appendix A Future Work
Appendix B Cost Analysis
In this work, we introduce FLIP. When using small language models under comparable inference budgets, FLIP achieves better performance than existing methods. The computational cost of FLIP is similar to that of LLM-as-a-Judge, as both approaches require a single inference call. However, the token budgets differ.
Specifically, we compare the input and output tokens required by each approach in Table 4. Since FLIP only takes the response as input, whereas LLM-as-a-Judge requires both the instruction and the response, the prompt length and consequently the inference cost is typically lower for FLIP under otherwise comparable settings.
Method Input Tokens Output Tokens
LLM-as-a-Judge instruction + response reasoning + answer
FLIP response reasoning + answer
Table 4: Comparison of token usage between FLIP and LLM-as-a-Judge.
Appendix C Robustness to Prompt Variations
@@
Table 5: Accuracy of the Llama 3 family models, averaged across the 1B, 3B, and 8B instruction-tuned variants, using various prompts on RewardBench2 across all subsets. “s.d.” denotes the standard deviation across different 3 prompts for each method. See Appendix I for the full list of prompts. FLIP is more robust to prompt variations than LLM-as-a-Judge.
Appendix D Similarity Metric
@@
Surprisingly, word-level F1 achieves better performance than embedding-based metrics, which are generally expected to better capture semantic similarity. We hypothesize that F1 surpasses BERTScore in this task for several reasons:
  • Sensitivity to related but over- or mis-specified instructions. In many cases, a response is not preferred because it addresses a related but different request, or because it over-extends beyond the intended scope. This often leads to an inferred instruction that is semantically similar but not exactly aligned with the original instruction. In such cases, F1 appropriately penalizes lexical differences, whereas BERTScore may still assign a high score due to embedding similarity. An example could be “List three causes of the French Revolution” and “Give an overview of the French Revolution, focusing on its causes”.
  • Instructions are short. Instructions and inferred instructions are typically short. In this regime, lexical overlap metrics such as F1 tend to be more stable, while embedding-based metrics may introduce noise and reduce discrimination.
  • Importance of key entities and exact wording. Instructions often contain critical nouns (e.g., names, locations, dates, or specific objects). If these tokens are missing or altered, the instruction meaning can change substantially. F1 captures these differences directly, while BERTScore may still give a high similarity score if the sentences remain broadly related. For example, “New York City” and “Los Angeles” have high embedding similarity but they are entirely different.
@@
Appendix E Generator Size
Appendix F Scaling Up
@@
Appendix G Test-Time Scaling with Parallel Sampling
Figures 8 and 9 present detailed results of test-time scaling with parallel sampling using the OLMo 2 and Llama 3 model families as reward models, respectively.
Appendix H Experiment Details
@@
Inference experiments are performed on either A100 or H200 GPUs. If a response does not follow the instructed output format or results in a tie, it will be considered incorrect and receive a score of 0.
@@
Appendix I Prompts
We list the prompts used in the main sets of experiments in Tables 9 and 10. Additional prompts used for the prompt consistency check are provided in Tables 11 through 14.
Appendix J Qualitative Examples
Table 9: Prompts used for all experiments, unless otherwise stated.
Table 10: Prompts used for all experiments, unless otherwise stated.